[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Oct 19, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Oct 19 19:18:32 UTC 2016


HSU Shift Log for Wed Oct 19 2016 15:01:12 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues: 12z guidance had the same hang-up issue that Dan described  
overnight. Mark helped guide us through the process at 12z and everything  
worked out fine. Monica fixed problem for 18Z guidance requests -- no  
issues.

N-AWIPS/AWIPS2 issues: None.

Other Techincal issues: None.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161019/cf1433d5/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list